# SGPMIL: Sparse Gaussian Process Multiple Instance Learning

### v1

@@ 3.2 SGPMIL: Sparse Gaussian Process MIL
where each $\mathbf{p_{s}}\in\Delta^{C-1}$ denotes the class probability vector for sample $s$ and $A_{s,k}$ is a matrix where each column contains $s$ samples for each instance $\mathbf{h_{k}}$ and for which $\frac{1}{S}\sum_{s}A_{s,k}\in\left[0,1\right]$. For $C$ classes, the $(C-1)$-dimensional probability simplex $\Delta^{C-1}=\left\{\mathbf{p}\in\mathbb{R}^{C}:p_{j}\geq 0\ \forall j\in[C],\ \sum_{j=1}^{C}p_{j}=1\right\}$. In turn, the joint probabilistic bag-level model can be written as,
@@
[…]
@@
In our approach, a learnable linear projection term in the variational posterior mean, defined as $\boldsymbol{\mu}_{\mathbf{X}}=W_{\mathbf{A}}X+b$. The parameters involved in the optimization include the weights of the projection layers and the classifier, collectively denoted by $W$; the projection matrix for the real inputs to the SGP $W_{\mathbf{A}}$; the inducing point set $\mathbf{Z}$; the kernel hyperparameters $\theta$ and the variational mean $\mathbf{m}_{\mathbf{U}}$. The kernel matrix elements are computed using a scaled Radial Basis Function (RBF) kernel function, defined as $\Sigma_{ij}=A\exp\left(-(x_{i}-x_{j})^{\top}\Theta^{-1}(x_{i}-x_{j})\right)+C$. These components are jointly learned by maximizing the Evidence Lower Bound (ELBO), which is defined as,
@@
[…]
3.3 Implementation Details
Our framework, encompassing data preprocessing and model training, is implemented in PyTorch [2] and PyTorch Lightning [9]. Feature extraction utilizes the TRIDENT toolkit and the UNI encoder [5]. For the AGP and SGPMIL models, we employ the AdamW optimizer [24] with a linear warmup followed by a cosine annealing learning rate schedule [23, 14], using a peak learning rate of $6\times 10^{-4}$ for the TCGA-NSCLC and CAMELYON16 datasets. All other models are trained using their original optimization settings. For the PANDA and BRACS datasets, all models are trained for a maximum of 20 epochs using the AdamW optimizer with cosine annealing and a maximum learning rate of $1\times 10^{-4}$. AGP and SGPMIL required a slightly higher peak learning rate of $2\times 10^{-4}$ to ensure convergence within the same number of epochs. The SGP layer is implemented using GPyTorch [11]. Complete code and instructions will be made publicly available to ensure full reproducibility.
@@ 4.1 Datasets
LM Post Diag $N_{U}$ ACC AUC ACE AUC ACE Time (s)
AGP [33] $\operatorname{softmax}(\cdot)$ $16$$858_{.029}.$$933_{.029}.$$089_{.007}.$$946$091.091.091.091 $31.0$/$6.0$
$80$$883_{.021}.$$954_{.019}.$$069_{.013}.$$953$080.080.080.080 $31.0$/$6.0$
$\sigma(\cdot)$ $80$$936_{.015}.$$970_{.009}.$$049_{.014}.$$822$320.320.320.320 $31.0$/$6.0$
$\operatorname{softmax}(\cdot)$ $80$$971_{.006}.$$976_{.007}.$$027_{.008}.$$821$173.173.173.173 $31.0$/$6.0$
$\sigma(\cdot)$ $80$$972_{.014}.$$979_{.016}.$$065_{.045}.$$970$103.103.103.103 $31.0$/$6.0$
$\operatorname{softmax}(\cdot)$ $80$$\underline{.973_{.007}}$$\underline{.980_{.009}}$$\underline{.024_{.007}}$$\underline{.965}$$\underline{.070}$ $\mathbf{9.0}$/$\mathbf{1.0}$
SGPMIL $\sigma(\cdot)$ $16$$971_{.016}.$$977_{.020}.$$045_{.022}.$$963$271.271.271.271 $\mathbf{9.0}$/$\mathbf{1.0}$
$80$$\mathbf{.980_{.007}}$$\mathbf{.986_{.005}}$$\mathbf{.021_{.005}}$$\mathbf{.973}$$\mathbf{.051}$ $\mathbf{9.0}$/$\mathbf{1.0}$
@@
Table 1: Ablation study on the CAMELYON16 dataset (10-fold cross-validation) evaluating architectural differences to vanilla AGP model. Variants differ in mean parameterization (LM), post-SGP activation (Softmax/Sigmoid), use of covariance diagonal (Diag) and number of inducing points ($N_{U}$). Metrics include WSI/Instance-level balanced accuracy (ACC), AUC and average training/inference time (T/P, in seconds).
CAMELYON16
@@
[8] includes 270 training and 130 test WSIs for breast cancer metastasis detection, with detailed tumor annotations for positive slides. We use 129 test slides due to a corrupted file. WSIs are segmented and patched at 10×10$\times$10 × magnification [26], using TRIDENT [43, 38], with features extracted via the UNI encoder [5]. We follow [25] for 10-fold cross-validation and report mean bag-level performance. Instance-level results use the model with the highest bag-level AUC.
TCGA-NSCLC
[…]
PANDA
@@
The PANDA dataset consists of 10,609 WSIs of prostate core needle biopsies, each annotated with an ISUP grade (0–5), resulting in a 6-class classification task. We perform 5-fold cross-validation using stratified splits, where each fold maintains approximately 80/5/15 splits per class for training, validation, and testing respectively. WSIs are segmented into non-overlapping $224\times 224$ patches at 20×20$\times$20 × magnification [36], and patch-level features are extracted using the UNI encoder [5, 36, 44].
BRACS
@@
The BRACS dataset [3] comprises 547 WSIs labeled as Benign, Atypical, or Malignant. We the predefined train, validation, and test split provided with the dataset, extract non-overlapping $224\times 224$ patches at 20×20$\times$20 × magnification [44], and obtain features using the UNI encoder [5]. All models are trained using 5 random seeds, and we report the mean and standard deviation of each evaluation metric.
4.2 Evaluation Metrics.
@@
For slide-level classification, we report balanced accuracy (ACC), area under the receiver operating characteristic curve (AUC), and adaptive expected calibration error (ACE) [31], capturing both predictive performance and model calibration.
@@ 5.1 Ablations
We showcase the performance gains achieved by the modifications introduced in our framework, relative to a baseline implementation of sparse Gaussian processes for MIL [33]. Table 1 showcases the improvement in bag-level performance our contributions have against the vanilla methodology. The ablations systematically vary four components: attention normalization (implemented via softmax or sigmoid), the mean parameterization in the variational posterior (LM), the use of diagonal covariance (Diag), and the number of inducing points ($N_{\mathbf{U}}$). The first improvement over the vanilla AGP model is achieved by introducing relaxed attention normalization, replacing the softmax constraint $\sum_{k}A_{s,k}=1$ with a sigmoid-based expectation constraint $\frac{1}{S}\sum_{s}A_{s,k}\in[0,1]$. This relaxation allows the model to naturally assign high attention to multiple instances with task-relevant characteristics, yielding initial gains in both bag- and instance-level performance.
Next, we observe a substantial improvement after introducing a feature-scaling term in the variational mean $\boldsymbol{\mu_{X}}=\mathbf{W_{A}}X+\mathbf{b}$, which replaces the constant $\mathbf{\mu_{X}}=C$ used in AGP. This adjustment improves expressiveness and makes optimization over kernel hyperparameters and inducing point locations $\mathbf{Z}$ more effective. The combination of relaxed normalization and feature scaling leads to further gains, especially in instance-level accuracy and calibration.
Finally, we address the numerical instability and speed bottlenecks of AGP by replacing the full covariance matrix with its diagonal approximation, eliminating the need for Cholesky decomposition during training. This leads to significantly faster and more stable training and inference, without sacrificing predictive performance.
Overall, introducing feature scaling in the variational mean, diagonal covariance approximation, and relaxed attention normalization lead to improved bag-level performance, faster convergence and more focused instance-level attention.
Figure 2: Slide-level results across multiple datasets.
@@ 5.2 Bag-Level Performance
Fig. 2 reports the performance of SGPMIL and several baselines across the different histopathology datasets and tasks. In particular, breast cancer detection (CAMELYON16), lung cancer subtyping (TCGA-NSCLC), prostate cancer grading (PANDA), and breast cancer subtyping (BRACS). Both PANDA and BRACS tasks are multiclass.
In CAMELYON16, all models—except vanilla AGP—achieve comparable accuracy and AUC. Notably, BayesMIL and SGPMIL demonstrate superior calibration, with BayesMIL attaining the lowest ACE, consistent with expectations from a probabilistic formulation. For TCGA-NSCLC, performance is uniformly high across models, indicating limited gains from architectural complexity for this binary task.
The PANDA dataset presents a greater challenge due to its multi-class nature. Here, our method ranks first in both ACC, AUC and $\kappa$, outperforming the next-best method by margins of $3.3\%$, $2.2\%$ and $2\%$, respectively, while ranking second in calibration. A similar trend is observed in BRACS, where our approach leads in accuracy and AUC by $3.3\%$ and $1.8\%$, again ranking second in calibration. We note that DGRMIL was not originally evaluated in multi-class settings. To integrate it, we substitute the binary cross-entropy loss with a cross-entropy loss followed by a softmax activation, updating the negative center when the bag belongs to the ’normal’ class and the positive center in any other case.
These results highlight the generalization capabilities of the SGP-based formulation, particularly in complex multi-class scenarios with heterogeneous tissue morphologies.
@@ 5.3 Instance-Level Localization
ABMIL [17] .$\ .769$$\ .885$$\ .216$$\ .517$$\ .552$$\ .057$
CLAM [25] .$\ .670$$\ .773$$\ .267$$\ .521$$\ .576$$\ \underline{.070}$
TransMIL [34] .$\ .598$$\ .840$$\ \underline{.063}$$\ .713$$\ \underline{.863}$$\ .217$
DGRMIL [45] .$\ .754$$\ .932$$\ .086$$\ .701$$\ .733$$\ .165$
BayesMIL [42] .$\ .806$$\ .862$$\ .182$$\ .576$$\ .643$$\ .120$
AGP [33] .$\ \underline{.863}$$\ \underline{.953}$$0$80\ .080 .$\ \underline{.738}$$\ .822$$\ .177$
SGPMIL .$\ \mathbf{.914}$$\ \mathbf{.973}$$\ \mathbf{.051}$$\ \mathbf{.768}$$\ \mathbf{.920}$$\ \mathbf{.044}$
Table 2: Instance-level performance on CAMELYON16 and BRACS for each model. Metrics shown: accuracy (ACC), AUC and ACE. Bold denotes best result.
Figure 3: Normalized attention heatmaps for all models. Scores are scaled to the interval $[0,1]$. For DGRMIL, we use the attention scores of the cls token, while for BayesMIL, AGP and SGPMIL we plot the mean attention score per patch. The yellow contours correspond to the ground-truth annotations
A key strength of our approach lies in its ability to identify relevant instances within each bag. We interpret the attention score as a proxy for the model’s capability to focus on patches with task-specific characteristics. For each dataset, we evaluate the best-performing model across all folds. For ABMIL and CLAM, attention scores are extracted from the gated attention mechanism. In TransMIL, we perform mean fusion of the attention matrices of attention heads within each of the two attention blocks and apply attention roll-out [1] to compute final scores. DGRMIL follows the method described by [45], using the global token attention. For BayesMIL, we perform multiple forward passes to sample attention scores and compute their mean per patch. For AGP and SGPMIL, we derive attention scores by averaging post-activation values—softmax for AGP and sigmoid for SGPMIL—across samples drawn from the respective learned SGP variational posteriors. All attention scores are subsequently normalized to the range $a\in[0,1]$ and evaluated against ground-truth annotations on test-set WSIs. In Table 2, SGPMIL consistently outperforms all baselines by over 3% in ACC and 2% in AUC, while providing more calibrated and focused attention on diagnostically salient regions.

### v2

@@ 3.2 SGPMIL: Sparse Gaussian Process MIL
where each $\mathbf{p_{s}}\in\Delta^{C-1}$ denotes the class probability vector for sample $s$ and $A_{s,k}$ is a matrix where each column contains $s$ samples for each instance $\mathbf{h_{k}}$ and for which $\frac{1}{S}\sum_{s}A_{s,k}\in\left[0,1\right]$. For $C$ classes, the (C1)(C-1)-dimensional probability simplex $\Delta^{C-1}=\left\{\mathbf{p}\in\mathbb{R}^{C}:p_{j}\geq 0\ \forall j\in[C],\ \sum_{j=1}^{C}p_{j}=1\right\}$. In turn, the joint probabilistic bag-level model can be written as,
@@
where $\mathbf{A}$ is the attention random variable, $\mathbf{U}$ are the inducing points and $\mathbf{Z}$, $\mathbf{X}$ are the learnable inducing point locations and actual data respectively. Following the work of [42], we place a prior over the inducing values $p(\mathbf{U})=\mathcal{N}\left(\boldsymbol{\mu}_{\mathbf{U}},\Sigma_{\mathbf{ZZ}}\right)$. To approximate the intractable joint distribution $p(\mathbf{A},\mathbf{U})=p(\mathbf{A}\mid\mathbf{U},\mathbf{Z},\mathbf{X})p(\mathbf{U}|\mathbf{Z})$ in Eq. 5, a variational distribution over the inducing values $q(\mathbf{U})=\mathcal{N}\left(\mathbf{U}\mid\mathbf{m}_{\mathbf{U}},\mathbf{S}_{\mathbf{ZZ}}\right)$ is considered. This leads to a closed form marginal for the attention distribution $q(\mathbf{A})=\int_{\mathbf{U}}p(\mathbf{A}\mid\mathbf{U})\,q(\mathbf{U})\,d\mathbf{U}$ which yields,
@@
In our approach, we introduce a learnable linear projection term in the variational posterior mean, defined as $\boldsymbol{\mu}_{\mathbf{X}}=W_{\mathbf{A}}X+b$. The parameters involved in the optimization include the weights of the projection layers and the classifier, collectively denoted by $W$; the projection matrix for the real inputs to the SGP $W_{\mathbf{A}}$; the inducing point set $\mathbf{Z}$; the kernel hyperparameters $\theta$ and the variational mean $\mathbf{m}_{\mathbf{U}}$. The kernel matrix elements are computed using a scaled Radial Basis Function (RBF) kernel function, defined as $\Sigma_{ij}=A\exp\left(-(x_{i}-x_{j})^{\top}\Theta^{-1}(x_{i}-x_{j})\right)+C$. These components are jointly learned by maximizing the Evidence Lower Bound (ELBO), which is defined as,
@@
where samples drawn from $q(\mathbf{A})$ are passed through the remainder of the network, producing samples of the class probability vectors $\mathbf{p}_{s}$ as defined in Eq. 4. Maximizing Eq. 8 is equivalent to minimizing a standard Cross-Entropy loss between predicted and observed class probabilities [14]. For inference, the predictive distribution is $p(\mathbf{Y^{*}})=\int_{\mathbf{A}}p(\mathbf{Y^{*}}\mid\mathbf{A})\,q(\mathbf{A})\,d\mathbf{A}$ and is approximated by taking MC samples from the learned variational posterior, propagating through the network and again computing the class probability vector as in Eq. 4. To generate MC samples of the attention scores, we apply the local reparameterization trick, rewriting $\mathbf{A}$ as $\mathbf{A}=\boldsymbol{m}_{\mathbf{A}}+\sqrt{\text{diag}(\mathbf{S}_{\mathbf{A}})}\odot\epsilon$ with $\mathbf{A}\sim q(\mathbf{A})$, where $\boldsymbol{m}_{\mathbf{A}}$ and $\mathbf{S}_{\mathbf{A}}$ represent the mean and covariance of the variational distribution respectively and $\epsilon\sim\mathcal{N}(0,I)$ the noise term.
@@ 4.1 Datasets
LM Post Diag $N_{U}$ ACC AUC ACE AUC ACE $t$ (s)
AGP [42] $\operatorname{softmax}(\cdot)$ $16$ $858_{.029} .$$933_{.029} .$$089_{.007} .$946.946 .091.091 31.031.0/$6.0$
$80$ $883_{.021} .$$954_{.019} .$$069_{.013} .$953.953 .080.080 31.031.0/$6.0$
$\sigma(\cdot)$ $80$ $936_{.015} .$$970_{.009} .$$049_{.014} .$822.822 .320.320 31.031.0/$6.0$
$\operatorname{softmax}(\cdot)$ $80$ $971_{.006} .$$976_{.007} .$$027_{.008} .$821.821 .173.173 31.031.0/$6.0$
$\sigma(\cdot)$ $80$ $972_{.014} .$$979_{.016} .$$065_{.045} .$$\underline{.970}$ .103.103 31.031.0/$6.0$
$\operatorname{softmax}(\cdot)$ $80$ $\underline{.973_{.007}}$ $\underline{.980_{.009}}$ $\underline{.024_{.007}}$ .965.965 $\underline{.070}$ $\mathbf{9.0}$$\mathbf{1.0}$
SGPMIL $\sigma(\cdot)$ $16$ $971_{.016} .$$977_{.020} .$$045_{.022} .$963.963 .271.271 $\mathbf{9.0}$$\mathbf{1.0}$
$80$ $\mathbf{.980_{.007}}$ $\mathbf{.986_{.005}}$ $\mathbf{.021_{.005}}$ $\mathbf{.973}$ $\mathbf{.051}$ $\mathbf{9.0}$$\mathbf{1.0}$
@@
Table 1: Ablation study on the CAMELYON16 dataset (10-fold cross-validation) evaluating architectural differences to vanilla AGP model. Variants differ in mean parameterization (LM), post-SGP activation (Softmax/Sigmoid), use of covariance diagonal (Diag) and number of inducing points ($N_{U}$). Metrics include WSI/Instance-level ACC, AUC, and average training/inference time $t$ (s).
@@
CAMELYON16 [11] includes 270 training and 130 test WSIs for breast cancer metastasis detection, with detailed tumor annotations for positive slides. We use 129 test slides due to a corrupted file. WSIs are segmented and patched at $10\times$ magnification [33], with features extracted via the UNI encoder [6]. We follow [32] for 10-fold cross-validation and report mean bag- and instance-level performance.
TCGA-NSCLC For the non-small cell lung carcinoma (NSCLC) subtyping task, we utilize hematoxylin and eosin (H&E) stained WSIs from The Cancer Genome Atlas (TCGA) to classify lung adenocarcinoma (LUAD) versus lung squamous cell carcinoma (LUSC) [7, 5]. The current cohort consists of 1,0061{,}006 slides, with $494$ LUAD and $512$ LUSC cases. WSIs are segmented and patched at $10{\times}$ magnification [33], with features extracted using UNI [6]. We evaluate performance using 4-fold cross-validation and report mean results across folds.
@@
PANDA [4] consists of 10,609 WSIs of prostate core needle biopsies, each annotated with an ISUP grade (0–5), resulting in a 6-class classification task. We perform 5-fold cross-validation using stratified splits, where each fold maintains approximately 80/5/15 splits per class for training, validation, and testing respectively. WSIs are segmented into non-overlapping $224\times 224$ patches at $20\times$ magnification [45], and features are extracted using UNI [6, 45, 58].
@@
BRACS [3] comprises 547 WSIs labeled as Benign, Atypical, or Malignant. We use the predefined train, validation, and test split provided with the dataset, extract non-overlapping $224\times 224$ patches at $20\times$ magnification [58], and obtain features using the UNI encoder [6]. All models are trained using 5 random seeds, and we report the mean and standard deviation of each evaluation metric.
4.2 Implementation Details
Our framework, encompassing data preprocessing and model training, is implemented in PyTorch [2] and PyTorch Lightning [13]. Feature extraction utilizes the TRIDENT toolkit [57, 49] and the UNI encoder [6]. For the AGP and SGPMIL models, we employ the AdamW optimizer [31] with a linear warmup followed by a cosine annealing learning rate schedule [30, 20], using a peak learning rate of $6\times 10^{-4}$ for the TCGA-NSCLC and CAMELYON16 datasets. All other models are trained using their original optimization settings. For the PANDA and BRACS datasets, all models are trained for a maximum of 20 epochs using the AdamW optimizer with cosine annealing and a maximum learning rate of $1\times 10^{-4}$. AGP and SGPMIL required a slightly higher peak learning rate of $2\times 10^{-4}$ to ensure convergence within the same number of epochs. The SGP layer is implemented using GPyTorch [16].
4.3 Evaluation Metrics
@@
For slide-level classification, we report balanced accuracy (ACC), area under the receiver operating characteristic curve (AUC), and adaptive expected calibration error (ACE) [39], capturing both prediction performance and model calibration. To evaluate statistical significance in all tasks, we conducted a one-sided paired t-test across cross-validation folds, testing whether our method consistently outperforms the baseline.
@@ 5.1 Ablations
We showcase the performance gains achieved by the modifications introduced in our framework, relative to a baseline implementation of sparse GPs for MIL [42]. Table 1 presents the improvement in bag-level performance our contributions have against the vanilla methodology. The ablations systematically vary four components: attention normalization (implemented via softmax or sigmoid), the mean parameterization in the variational posterior (LM), the use of diagonal covariance (Diag), and the number of inducing points ($N_{\mathbf{U}}$). Analysis on the effect of varying the number of inducing points is provided in supplementary Table 11.
The first improvement over the vanilla AGP model is achieved by introducing relaxed attention normalization, replacing the softmax constraint $\sum_{k}A_{s,k}=1$ with a sigmoid-based expectation constraint $\frac{1}{S}\sum_{s}A_{s,k}\in[0,1]$. This relaxation allows the model to naturally assign high attention to multiple instances with task-relevant characteristics, yielding performance gains in both bag- and instance-level. Next, we observe a substantial improvement after introducing a feature-scaling term in the variational mean $\boldsymbol{\mu_{X}}=\mathbf{W_{A}}X+\mathbf{b}$, which replaces the constant $\mathbf{\mu_{X}}=C$ used in AGP. This adjustment improves expressiveness and makes optimization over kernel hyperparameters and inducing point locations $\mathbf{Z}$ more effective. The combination of relaxed normalization and feature scaling leads to further gains, especially in instance-level accuracy and calibration.
Finally, we address the numerical instability and speed bottlenecks of AGP by replacing the full covariance matrix with its diagonal approximation, eliminating the need for Cholesky decomposition during training. This leads to much faster and more stable training and inference, without sacrificing predictive performance. Comprehensive training and inference time comparisons against all baselines are provided in supplementary Table 6. To further underscore the stability gains of our modifications, we compute the Lipschitz constant as a surrogate of stability of each variant $L(f)\;\leq\;\prod_{\ell=1}^{L}\,\sigma_{\max}(W_{\ell})\,\prod_{\ell=1}^{L-1}L_{\sigma_{\ell}}$, where $\sigma_{max}$ denotes the largest singular value, and $L_{\sigma_{l}}$ the Lipschitz value of activation functions, obtaining $L_{\mathrm{AGP}}=105.5_{25.9}$ vs. $L_{\mathrm{SGPMIL}}=38.1_{9.1}$, a major reduction in input sensitivity, underscoring the significance of our contributions. Overall, introducing feature scaling in the variational mean, diagonal covariance approximation, and relaxed attention normalization leads to improved bag-level performance, faster convergence, and more focused instance-level attention.
Figure 2: Slide-level performance across multiple datasets. denotes statistical significance ($p<0.05$) based on one-sided paired t-tests.
@@ 5.2 Bag-Level Performance
Figure 2 reports the performance of SGPMIL and several baselines across the different histopathology datasets and tasks together with one-sided paired t-tests to evaluate significance for all datasets.
On CAMELYON16, SGPMIL achieves significantly higher accuracy than all models except CLAM and BayesMIL, and significantly higher AUC than all models except ABMIL and CLAM. In terms of calibration, BayesMIL, CLAM, and SGPMIL perform comparably well, with SGPMIL achieving statistically significantly lower ACE than the remaining approaches, while BayesMIL attains the lowest ACE overall, consistent with its probabilistic formulation. For TCGA-NSCLC, performance is uniformly high across models; however, SGPMIL is significantly better than CLAM in ACC and AUC, outperforms MixMIL and AGP in ACC and achieves better calibration than MixMIL.
The PANDA dataset presents a greater challenge due to its multi-class nature. Here, our method ranks first with statistically significant improvements in both ACC and $\kappa$—the preferred metric over AUC for this task—outperforming the next-best method by margins of $3.3\%$, and $2.2\%$ respectively, while achieving comparable calibration to AGP and superior calibration to the other baselines. A similar trend is observed in BRACS, where our approach leads in both ACC and AUC by $3.3\%$ and $1.5\%$ in a significant manner, again performing similarly to AGP in calibration and better than other baselines. We note that DGRMIL was not originally evaluated in multi-class settings. To integrate it, we substitute the binary cross-entropy loss with a cross-entropy loss followed by a softmax activation, updating the negative center when the bag belongs to the ’normal’ class and the positive center in any other case. These results highlight the generalization capabilities of the SGP-based formulation, particularly in complex multi-class scenarios with heterogeneous tissue morphologies.
@@ 5.3 Instance-Level Localization
ABMIL [23] $775^{\star}\$$910^{\star}\$$191^{\star}\$$514^{\star}\$$560^{\star}\$.062\ .062\
[…]
MixMIL [12] $563^{\star}\$$670^{\star}\$$256^{\star}\$$509^{\star}\$$528^{\star}\$.033\ $\mathbf${.033}\
[…]
SGPMIL .910\ $\mathbf${.910}\ .979\ $\mathbf${.979}\ .059¯\ $\underline${.059}\ .765\ $\mathbf${.765}\ .899\ $\mathbf${.899}\ .050¯\ $\underline${.050}\
Table 2: Instance-level performance on CAMELYON16 and BRACS for each model. Metrics shown: accuracy (ACC), AUC and ACE. Bold denotes best result. denotes statistical significance ($p<0.05$) based on one-sided paired t-tests.
Figure 3: Inducing-point structure and normalized attention heatmaps on a CAMELYON16 test slide. Top row: (a) slide with ground-truth annotations; (b) inducing-point label map; (c) top-7 most similar patches for the most representative inducing points. Bottom row: normalized attention heatmaps (scores in $[0,1]$). For DGRMIL, we use the cls-token attention; for BayesMIL, AGP, and SGPMIL, we plot the mean attention per patch. Yellow contours denote ground-truth annotations. See Supplementary Figure 3 for additional visualizations.
A key strength of our approach lies in its ability to identify relevant instances within each bag. We interpret the attention score as a proxy for the model’s capability to focus on patches with task-specific characteristics. For each dataset, we report the performance across all folds. For ABMIL and CLAM, attention scores are extracted from the gated attention mechanism. In TransMIL, we perform mean fusion of the attention matrices of attention heads within each of the two attention blocks and apply attention roll-out [1] to compute final scores. DGRMIL follows the method described by [59], using the global token attention. For BayesMIL and MixMIL, we use the mean of the attention variational posterior per patch. For AGP and SGPMIL, we derive attention scores by averaging post-activation values—softmax for AGP and sigmoid for SGPMIL—across samples drawn from the learned SGP variational posteriors. All attention scores are subsequently normalized to the range $a\in[0,1]$ and evaluated against ground-truth annotations on test-set WSIs. In Table 2, SGPMIL consistently outperforms all baselines by over 4% in ACC and 2% in AUC in a statistically significant manner; calibration-wise AGP and SGPMIL are more calibrated than all other baselines in CAMELYON16 while on BRACS, SGPMIL achieves the best calibration among all methods with non-trivial performance, as ABMIL, CLAM and MixMIL perform at chance level. Moreover, SGPMIL provides focused attention on diagnostically salient regions.